# Vibravox: A Dataset of French Speech Captured with Body-conduction Audio Sensors

### v3

@@ Abstract
Vibravox is a dataset compliant with the General Data Protection Regulation (GDPR) containing audio recordings using five different body-conduction audio sensors : two in-ear microphones, two bone conduction vibration pickups and a laryngophone. The dataset also includes audio data from an airborne microphone used as a reference. The Vibravox corpus contains 45 hours of speech samples and physiological sounds recorded by 188 participants under different acoustic conditions imposed by an high order ambisonics 3D spatializer. Annotations about the recording conditions and linguistic transcriptions are also included in the corpus. We conducted a series of experiments on various speech-related tasks, including speech recognition, speech enhancement and speaker verification. These experiments were carried out using state-of-the-art models to evaluate and compare their performances on signals captured by the different audio sensors offered by the Vibravox dataset, with the aim of gaining a better grasp of their individual characteristics.
@@ I Introduction
Research into physics and electronics is improving with some skin-attachable sensors such as [6, 7, 8]. Similarly to earlier bone and throat microphones, these new wearable sensors detect skin vibration, which is highly and linearly correlated with the acoustic pressure produced by the wearer’s voice [9]. Not only do they improve on the state of the art by having superior sensitivity over the vocal frequency range – thereby improving the signal-to-noise ratio – but they also have superior skin compliance, which facilitates adhesion to curved skin surfaces. These new kinds of sensors, just like the previous ones, are however unable to capture the full bandwidth of the speech signal, due to the inherent low-pass filtering of tissues. The manufacturing process also needs to be stabilized, which currently prevents them from being commercially available. In addition, some earbuds, such as the Xiaomi Buds 4 Pro, already include some voice pickup unit thanks to a retro-action microphone integrated into the loudspeaker at the entrance to the ear canal. This exemplifies the significant role that BCMs will likely assume in the future.
@@
Deep learning methods have shown excellent performance in a variety of speech and audio tasks. In this work, we show that the Vibravox dataset can be used to advance research in three key tasks that would enable overcoming the current limitations of body conduction sensors: bandwidth extension, speech recognition, and speaker verification. Since body conduction sensors exhibit reduced bandwidth transduction efficiency, works such as [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] demonstrate the ability of deep learning approaches to regenerate mid and high frequencies from low frequency audio content. They have adapted recent deep learning trends to the specific problem of bandwidth extension, also known as audio super-resolution. For robust speech recognition, models such as Whisper [29] or Canary1B [30] have pushed the limits of usable signals. Finally, for speaker verification, approaches such as TitaNet [31], WavLM [32], Pyannote [33] and ECAPA2 [34] now allow a wide range of signals to be used thanks to their increased robustness.
@@
The availability of large datasets is critical to advancing research and development of BCMs. These datasets allowing to train and evaluate deep learning models have been a key missing ingredient in achieving high-quality, intelligible speech with such audio sensors. The assumption that airborne and body-transmitted speech share identical excitation sources and possess a simple transfer function between them being inadequate, producing realistic synthetic data is therefore still a challenging task. In a previous study we performed Out of Distribution tests performed on real BCM signals with an EBEN model trained on simulated data [17], showing the importance of using a sufficiently large set of real data for training. Despite their importance, existing datasets still exhibit several notable limitations. Data collection being labor-intensive, the existing BCM datasets often lack the necessary scale and diversity to comprehensively cover the full range of acoustic scenarios encountered in real-world applications. Another limitation is the absence of publicly available BCM datasets in French, restricting research primarily to the languages covered by existing resources. Problems also persist with signal quality, including noise and artifacts. On top of these issues, the variety of sensors used in existing datasets remains limited, making it difficult to generalize results across different recording environments and equipment configurations. The Vibravox dataset is being made available to fill these gaps and stimulate research in the field of speech capture using non-conventional audio sensors.
@@
Dataset Number of speakers Number of BCM Clean speech Noisy speech Data augmentation Language Text transcript Phonetic transcript Download page
@@
In this paper, we contribute along two major axes: the construction and utilization of the dataset. First, we introduce the Vibravox dataset, developed through a rigorous data collection process. We designed and implemented a recording protocol, assembled the necessary hardware, and built dedicated software to ensure a controlled and high-quality data acquisition setup. A total of 200 participants contributed to the dataset, with 188 retained after post-processing to ensure high-quality data. We applied post-filtering techniques to enhance data quality and facilitate interpretability, enabling meaningful comparisons between different microphones and insightful speech signal analyses. Second, we establish baselines using the Vibravox dataset for three key speech processing tasks: bandwidth extension, speech-to-phoneme recognition, and speaker verification. For each task, we conduct experiments in both quiet and noisy environments to comprehensively evaluate the strengths and limitations of body-conduction microphones under realistic conditions. For speech-to-phoneme recognition and speaker verification, we compare the performance of raw and enhanced body-conduction sensors in quiet conditions to evaluate the impact of speech enhancement on downstream applications. These contributions provide new insights into the potential and challenges of BCMs for speech capture and processing, paving the way for future advancements in the field.
@@ II-A Training datasets
Before delving into publicly available datasets, it is worth acknowledging that they have been made available for a relatively short period of time, are in some cases of insufficient size, and are only available in a restricted set of languages. Consequently, several studies have employed low-pass filtering of high-quality audio to artificially generate body-conducted speech data. A hybrid approach involves the collection of individual transfer functions between a reference microphone and a body-conducted microphone, which is the focus of the Hearpiece database, as referenced in [40]. In their article, M. Ohlenbusch et al. [41] proposed several speech-dependent models of one’s own voice transfer function to simulate the degradation induced by body conducted audio recordings. Their results indicate that the transfer function is of course speaker-dependent, but also phoneme-dependent due to the different position of the jaw. Their study brings a significant improvement in the ability to simulate data. Following on from this work, the authors also published [42] to propose speech-dependent data augmentation to compensate for the lack of a large dataset of own speech signals. They were able to reduce the performance gap when testing the model on real signals. However, by analogy with FineWeb-Edu [43], the performance improvement can also come from having a larger and highly curated dataset on which to perform classical supervised training.
@@
A comprehensive literature review of existing, publicly available datasets of BCM-captured speech is essential to gain a full understanding of the subject. This encompasses all forms of capture, including bone, in-ear, and throat sensors. It should be noted that there exists a number of small private datasets [44, 45, 46, 47, 48, 49], albeit not open-sourced. The few remaining public datasets are listed in Table I, including Vibravox. In terms of size, the ESMB dataset [36] is the largest, with 128 hours of recorded speech, followed by Vibravox with its 45h37 of recordings, ABCS [35] (42 hours), and EmoBone [37] (19 hours). The final dataset to be considered is TAPS (13h18) [38]. This dataset includes post-recording processing to realign signals at the sample level, thereby easing the direct utilization of L1 loss in the wavform domain for the training of deep learning models to enhance body-conducted speech. Although Vibravox is not the largest dataset, it overcomes several limitations found in existing datasets, such as the limited diversity of audio sensors and the lack of noise recordings. Vibravox addresses these issues by using five different BCMs and including recordings of speech in both noisy and quiet environments. Previous studies indicate that the transmission of external noise through the device is not only influenced by the device itself, but also by individual variance [40] and angle of arrival [50]. For this reason, the Vibravox dataset was recorded in a 3D sound spatializer to sample uniformly the sphere for noise emission. Its unique combination of simultaneous multi-microphone recordings and French-language data therefore provides complementary resources to existing datasets, offering new perspectives for non conventional speech capture and comparative analyses of different sensor technologies.
@@ II-B Speech processing tasks
The body-conducted speech research community has made significant progress in recent years. The accessibility of public datasets such as Vibravox has become increasingly important for the development and improvement of deep learning models in this field. The continued availability of these resources is likely to facilitate further advances and innovations in body-conducted speech technology, for tasks such as speech enhancement, speech transcription, and speaker verification.
@@ II-B1 Speech Enhancement
BCM datasets, paired with airborne speech, are invaluable in the field of speech enhancement. This task refers to the process of improving the quality and intelligibility of speech, often in the presence of background noise. However, speech enhancement on BCM is mainly about bandwidth extension, as the sensors are robust to noise but cannot record high frequencies due to physical constraints. Several approaches have been published recently in this context. Among them, [49, 20, 25] model the long-term dependencies of speech by adapting the non-quadratic complexity variations of the attention mechanism [51, 52, 53]. Some studies show superior reconstruction metrics compared to EBEN, but as shown in Table 1 of [16], one of the approaches reporting the best metrics is Kuleshov’s U-net [54], which is also the one that performs worst in the MUSHRA (MUltiple Stimuli with Hidden Reference and Anchor) analysis. This evidence serves to illustrate the importance of conducting listening tests when comparing reconstructive and generative approaches. Furthermore, EBEN’s computational and memory requirements are low, due not only to the limited number of parameters in the generator, but also to the highly strided embeddings facilitated by the Pseudo-Quadrature Mirror Filters (PQMF) representation [17]. In [23], the authors proposed an innovative approach to deal with limited datasets. They developed a method to disentangle a speaker embedding vector from the hidden representations of the network during the enhancement process, which allowed a fair generalization performance with a limited training set of 5 hours of 29 speakers and additional synthetic data. Conversely, the approach proposed in [25] involves pre-training with simulated data and fine-tuning on a device with real data, thereby creating a speaker-dependent speech enhancement model.
@@ II-B2 Speech-to-text
Transcribed speech datasets are essential for training models to accurately recognize and transcribe spoken words. Eventually, performances can be boosted with a self-supervised pretraining [55] on unlabeled speech data to construct a semantic representation of speech prior to using text labels. To perform robust transcription, the inclusion of speech in both noisy and quiet environments in the Vibravox dataset is particularly beneficial, as it allows for the development of models that can perform well in a variety of acoustic conditions. To provide context, [56, 35] have previously addressed this task using this particular data modality. They have shown that the body-conducted speech can improve the performance of a basic speech recognition system in adverse environments solely based on airborne speech.
@@ II-B3 Speaker verification
Finally, speaker verification is a biometric authentication method that uses a person’s voice to confirm their identity. As with all the other tasks, if the microphone is less affected by reverberation and external noise, the problem of domain mismatch is reduced. This, in turn, enables higher recognition accuracy in challenging environments. Another benefit that is not immediately apparent is the amount of information available. Because everyone’s skull and timbre are unique, the recorded signal can be more descriptive of a person’s identity. This wealth of biometric information opens up several potential applications, including those in the military domain, as detailed in [57]. One such application is access control, where it can strengthen security by adding an extra layer compared to conventional methods like keycards or codes. Another compelling application lies in intelligence gathering. By pinpointing known voices in intercepted enemy transmissions, speaker identification can reveal the involvement of high-value targets. Beyond this paper’s focus, this rich biometric data (heart rate, breathing) could be a game-changer in battlefield forensics, aiding revealing emotional states and tracking life signal during critical events.
@@ III-A Hardware
Forehead accelerometer Forehead One-axis piezeoceramic accelerometer Knowles BU23173-000 23.7 dB -3.5 dB
@@
A comprehensive list of all BCMs employed in the Vibravox dataset is presented in Tab. II. Their positioning on the participants’ heads is shown in Figure 1. Although not depicted in Figure 1, a custom 3D-printed backpack and headset were also designed with the objective of facilitating the optimal placement of sensors, carrying the Zoom F8n field recorder, and routing audio cables to that device111several pictures and a video of participants wearing the sensors and the backpack during experiments are available at https://vibravox.cnam.fr/documentation/hardware/sensors. The Zoom F8n records all six sensor signals simultaneously, operates at a sampling rate of 48 kHz and a resolution of 32 bits, thereby guaranteeing precise signal capture with high fidelity. To enhance audio data quality, a $20$ Hz high-pass filter is uniformly applied to each acquisition channel, effectively attenuating low-frequency noise and artifacts.
@@
In order to provide a balanced comparison between body conduction and airborne audio sensors in terms of immunity to ambient noise, we selected an airborne microphone mounted on a headset that presents a cardioid directivity pattern towards the speaker’s mouth, hence producing a dry reference. The in-ear soft foam-embedded microphone is a prototype designed by the ISL and used in [58]. The in-ear rigid earpiece-embedded microphone was initially presented in [59] as a versatile research earpiece with multiple drivers and microphones. In this study, we use the occluded version of this device, and only record the in-ear microphone signal. The laryngophone has been purchased from iXradio. This device consists of dual piezoelectric sensors on an adjustable neck rim, which captures skin vibrations from vocal cords and targets voice amplifiers, recorders, and Voice over Internet Protocol (VoIP) applications. The forehead accelerometer, affixed via a homemade headband, is a compact contact sensor that has been commercialized by Knowles. It is designed for use in high-noise environments and can be integrated into helmets for military and emergency service communications. Finally, the temple vibration sensor is an AKG product, a miniature vibration sensor originally designed for string instruments. In this study, this sensor has been diverted from its original purpose to capture bone vibrations by placing it in an adjustable 3D-printed enclosure adapted to the participants’ zygomatic bones. In addition to sensor specifications, Table II presents two signal-to-noise ratio (SNR) measures for each sensor. The first, signal-to-self-noise, is computed as the difference between the $L_{50}^{silence}$ and $L_{50}^{speech}$ values, which will be defined in Section III-E, across the entire speechless-clean split of the dataset, as defined in Section III-D. Notably, all sensors exhibit a high signal-to-self-noise ratio, except for the temple vibration sensor, which is primarily designed to capture much higher amplitude signals. The second measure is the signal-to-external-noise ratio, calculated under an ambient noise level of 90 dB. Since the reference signal is unavailable in this case, we derived this ratio by decomposing it into the signal-to-self-noise minus the external-noise-to-self-noise ratio. This was estimated using the speechless-noisy and speechless-clean subsets, as defined in Section III-D. This decomposition assumes that speakers produce speech at the same loudness in both quiet and noisy environments. However, in practice, the Lombard effect causes speakers to increase their vocal effort in noisy conditions, meaning the reported signal-to-external-noise ratios represent a lower bound on the actual values. Despite this approximation, the signal-to-external-noise ratio remains a direct indicator of each sensor’s noise resilience. Our results show that sensors directly capturing skin vibrations exhibit the highest noise resilience. However, these sensors also tend to have the smallest bandwidth, highlighting the trade-off between noise robustness and frequency range.

### v4

@@ Abstract
Vibravox is a dataset compliant with the General Data Protection Regulation (GDPR) containing audio recordings using five different body-conduction audio sensors: two in-ear microphones, two bone conduction vibration pickups, and a laryngophone. The dataset also includes audio data from an airborne microphone used as a reference. The Vibravox corpus contains 45 hours per sensor of speech samples and physiological sounds recorded by 188 participants under different acoustic conditions imposed by a high order ambisonics 3D spatializer. Annotations about the recording conditions and linguistic transcriptions are also included in the corpus. We conducted a series of experiments on various speech-related tasks, including speech recognition, speech enhancement, and speaker verification. These experiments were carried out using state-of-the-art models to evaluate and compare their performances on signals captured by the different audio sensors offered by the Vibravox dataset, with the aim of gaining a better grasp of their individual characteristics.
@@ I Introduction
Research into physics and electronics is improving with some skin-attachable sensors such as [6, 7, 8]. Similarly to earlier bone and throat microphones, these new wearable sensors detect skin vibration, which is highly and linearly correlated with the acoustic pressure produced by the wearer’s voice [9]. Not only do they improve on the state of the art by having superior sensitivity over the vocal frequency range – thereby improving the signal-to-noise ratio – but they also have superior skin compliance, which facilitates adhesion to curved skin surfaces. These new kinds of sensors, just like the previous ones, are, however, unable to capture the full bandwidth of the speech signal due to the inherent low-pass filtering of tissues. The manufacturing process also needs to be stabilized, which currently prevents them from being commercially available. In addition, some earbuds, such as the Xiaomi Buds 4 Pro, already include some voice pickup unit thanks to a retro-action microphone integrated into the loudspeaker at the entrance to the ear canal. Finally, the innovative Aria Gen 2 Meta’s research platform now integrates a contact microphone. This exemplifies the significant role that BCMs will likely assume in the future.
@@
Deep learning methods have shown excellent performance in a variety of speech and audio tasks. In this work, we show that the Vibravox dataset can be used to advance research in three key tasks that would enable overcoming the current limitations of body conduction sensors: bandwidth extension, speech recognition, and speaker verification. Since body conduction sensors exhibit reduced bandwidth transduction efficiency, works such as [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] demonstrate the ability of deep learning approaches to regenerate mid and high frequencies from low-frequency audio content. They have adapted recent deep learning trends to the specific problem of bandwidth extension, also known as audio super-resolution. For robust speech recognition, models such as Whisper [29] or Canary1B [30] have pushed the limits of usable signals. Finally, for speaker verification, approaches such as TitaNet [31], WavLM [32], Pyannote [33], and ECAPA2 [34] now allow a wide range of signals to be used thanks to their increased robustness.
@@
The availability of large datasets is critical to advancing research and development of BCMs. These datasets, allowing the training and evaluation of deep learning models, have been a key missing ingredient in achieving high-quality, intelligible speech with such audio sensors. The assumption that airborne and body-transmitted speech share identical excitation sources and possess a simple transfer function between them is inadequate; producing realistic synthetic data is therefore still a challenging task. In a previous study, we performed out-of-distribution tests performed on real BCM signals with an EBEN model trained on simulated data [17], showing the importance of using a sufficiently large set of real data for training. Despite their importance, existing datasets still exhibit several notable limitations. Data collection being labor-intensive, the existing BCM datasets often lack the necessary scale and diversity to comprehensively cover the full range of acoustic scenarios encountered in real-world applications. Another limitation is the absence of publicly available BCM datasets in French, restricting research primarily to the languages covered by existing resources. Problems also persist with signal quality, including noise and artifacts. On top of these issues, the variety of sensors used in existing datasets remains limited, making it difficult to generalize results across different recording environments and equipment configurations. The Vibravox dataset is being made available to fill these gaps and stimulate research in the field of speech capture using non-conventional audio sensors.
@@
Dataset Number of speakers Number of BCM Clean speech Noisy speech Speechless recordings Language Text transcript Phonetic transcript Download page
@@
In this paper, we contribute along two major axes: the construction and utilization of the dataset. First, we introduce the Vibravox dataset, developed through a rigorous data collection process. We designed and implemented a recording protocol, assembled the necessary hardware, and built dedicated software to ensure a controlled and high-quality data acquisition setup. A total of 200 participants contributed to the dataset, with 188 retained after post-processing to ensure high-quality data. We applied post-filtering techniques to enhance data quality and facilitate interpretability, enabling meaningful comparisons between different microphones and insightful speech signal analyses. Second, we establish baselines using the Vibravox dataset for three key speech processing tasks: bandwidth extension, speech-to-phonemes recognition, and speaker verification. For each task, we conduct experiments in both quiet and noisy environments to comprehensively evaluate the strengths and limitations of body-conduction microphones under realistic conditions. For speech-to-phonemes recognition and speaker verification, we compare the performance of raw and enhanced body-conduction sensors in quiet conditions to evaluate the impact of speech enhancement on downstream applications. These contributions provide new insights into the potential and challenges of BCMs for speech capture and processing, paving the way for future advancements in the field.
@@ II-A Training datasets
Before delving into publicly available datasets, it is worth acknowledging that they have been made available for a relatively short period of time, are in some cases of insufficient size, and are only available in a restricted set of languages. Consequently, several studies have employed low-pass filtering of high-quality audio to artificially generate body-conducted speech data. A hybrid approach involves the collection of individual transfer functions between a reference microphone and a body-conducted microphone, which is the focus of the Hearpiece database, as referenced in [40]. In their article, Ohlenbusch et al. [41] proposed several speech-dependent models of one’s own voice transfer function to simulate the degradation induced by body-conducted audio recordings. Their results indicate that the transfer function is, of course, speaker-dependent, but also phoneme-dependent due to the different position of the jaw. Their study brings a significant improvement in the ability to simulate data. Following on from this work, the authors also published [42] to propose speech-dependent data augmentation to compensate for the lack of a large dataset of own speech signals. They were able to reduce the performance gap when testing the model on real signals. However, by analogy with FineWeb-Edu [43], the performance improvement can also come from having a larger and highly curated dataset on which to perform classical supervised training.
@@
A comprehensive literature review of existing, publicly available datasets of BCM-captured speech is essential to gain a full understanding of the subject. This encompasses all forms of capture, including bone, in-ear, and throat sensors. It should be noted that there exist several small private datasets [44, 45, 46, 47, 48, 49, 50], albeit not open-sourced. The few remaining public datasets are listed in Table I, including Vibravox. In terms of size, the ESMB dataset [36] is the largest, with 128 hours of recorded speech, followed by Vibravox with its 45h37 of recordings per sensor, ABCS [35] (42 hours), and EmoBone [37] (19 hours). The final dataset to be considered is TAPS (13h18) [38]. This dataset includes post-recording processing to realign signals at the sample level, thereby easing the direct utilization of L1 loss in the waveform domain for the training of deep learning models to enhance body-conducted speech. Although Vibravox is not the largest dataset, it overcomes several limitations found in existing datasets, such as the limited diversity of audio sensors and the lack of noise recordings. Vibravox addresses these issues by using five different BCMs and including recordings of speech in both noisy and quiet environments. Previous studies indicate that the transmission of external noise through the device is not only influenced by the device itself but also by individual variance [40] and angle of arrival [51]. For this reason, the Vibravox dataset was recorded in a 3D sound spatializer to sample uniformly the sphere for noise emission. Its unique combination of simultaneous multi-microphone recordings and French-language data therefore provides complementary resources to existing datasets, offering new perspectives for non-conventional speech capture and comparative analyses of different sensor technologies.
@@ II-B Speech processing tasks
The body-conducted speech research community has made significant progress in recent years. The accessibility of public datasets such as Vibravox has become increasingly important for the development and improvement of deep learning models in this field. The continued availability of these resources is likely to facilitate further advances and innovations in body-conducted speech technology for tasks such as speech enhancement, speech transcription, and speaker verification.
@@ II-B1 Speech Enhancement
BCM datasets, paired with airborne speech, are invaluable in the field of speech enhancement. This task refers to the process of improving the quality and intelligibility of speech, often in the presence of background noise. However, speech enhancement on BCM is mainly about bandwidth extension, as the sensors are robust to noise but cannot record high frequencies due to physical constraints. Several approaches have been published recently in this context. Among them, [49, 20, 25] model the long-term dependencies of speech by adapting the non-quadratic complexity variations of the attention mechanism [52, 53, 54]. Some studies show superior reconstruction metrics compared to EBEN, but as shown in Table 1 of [16], one of the approaches reporting the best metrics is Kuleshov’s U-net [55], which is also the one that performs worst in the MUSHRA (MUltiple Stimuli with Hidden Reference and Anchor) analysis. This evidence serves to illustrate the importance of conducting listening tests when comparing reconstructive and generative approaches. Furthermore, EBEN’s computational and memory requirements are low, due not only to the limited number of parameters in the generator but also to the highly strided embeddings facilitated by the Pseudo-Quadrature Mirror Filters (PQMF) representation [17]. In [23], the authors proposed an innovative approach to deal with limited datasets. They developed a method to disentangle a speaker embedding vector from the hidden representations of the network during the enhancement process, which allowed a fair generalization performance with a limited training set of 5 hours of 29 speakers and additional synthetic data. Conversely, the approach proposed in [25] involves pre-training with simulated data and fine-tuning on a device with real data, thereby creating a speaker-dependent speech enhancement model.
@@ II-B2 Speech-to-text
Transcribed speech datasets are essential for training models to accurately recognize and transcribe spoken words. Eventually, performances can be boosted with a self-supervised pretraining [56] on unlabeled speech data to construct a semantic representation of speech prior to using text labels. To perform robust transcription, the inclusion of speech in both noisy and quiet environments in the Vibravox dataset is particularly beneficial, as it allows for the development of models that can perform well in a variety of acoustic conditions. To provide context, [57, 35, 58] have previously addressed this task using this particular data modality. They have shown that body-conducted speech can improve the performance of a basic speech recognition system in adverse environments solely based on airborne speech.
@@ II-B3 Speaker verification
Finally, speaker verification is a biometric authentication method that uses a person’s voice to confirm their identity. As with all the other tasks, if the microphone is less affected by reverberation and external noise, the problem of domain mismatch is reduced. This, in turn, enables higher recognition accuracy in challenging environments. Another benefit that is not immediately apparent is the amount of information available. Because everyone’s skull and timbre are unique, the recorded signal can be more descriptive of a person’s identity. This wealth of biometric information opens up several potential applications, including those in the military domain, as detailed in [59]. One such application is access control, where it can strengthen security by adding an extra layer compared to conventional methods like keycards or codes. Another compelling application lies in intelligence gathering. By pinpointing known voices in intercepted transmissions, speaker identification can reveal the involvement of high-value targets. Beyond this paper’s focus, this rich biometric data (heart rate, breathing) could be a game-changer in battlefield forensics, aiding in revealing emotional states and tracking life signals during critical events.
@@ III Building Vibravox
Forehead accelerometer Forehead One-axis piezoceramic accelerometer Knowles BU23173-000 23.7 dB -3.5 dB
@@ III-A Hardware
A comprehensive list of all BCMs employed in the Vibravox dataset is presented in Tab. II. Their positioning on the participants’ heads is shown in Figure 1. Although not depicted in Figure 1, a custom 3D-printed backpack and headset were also designed with the objective of facilitating the optimal placement of sensors, carrying the Zoom F8n field recorder, and routing audio cables to that device [^1] Several pictures and a video of participants wearing the sensors and the backpack during experiments are available at https://vibravox.cnam.fr/documentation/hardware/sensors. The Zoom F8n records all six sensor signals simultaneously, operates at a sampling rate of 48 kHz and a resolution of 32 bits, thereby guaranteeing precise signal capture with high fidelity. To enhance audio data quality, a 20 Hz high-pass filter is uniformly applied to each acquisition channel, effectively attenuating low-frequency noise and artifacts.
@@
In order to provide a balanced comparison between body conduction and airborne audio sensors in terms of immunity to ambient noise, we selected an airborne microphone mounted on a headset that presents a cardioid directivity pattern towards the speaker’s mouth, hence producing a dry reference. The in-ear soft foam-embedded microphone is a prototype designed by the ISL and used in [60]. The in-ear rigid earpiece-embedded microphone was initially presented in [61] as a versatile research earpiece with multiple drivers and microphones. In this study, we use the occluded version of this device and only record the in-ear microphone signal. The laryngophone has been purchased from iXradio. This device consists of dual piezoelectric sensors on an adjustable neck rim, which captures skin vibrations from vocal cords and targets voice amplifiers, recorders, and Voice over Internet Protocol (VoIP) applications. The forehead accelerometer, affixed via a homemade headband, is a compact contact sensor that has been commercialized by Knowles. It is designed for use in high-noise environments and can be integrated into helmets for military and emergency service communications. Finally, the temple vibration sensor is an AKG product, a miniature vibration sensor originally designed for string instruments. In this study, this sensor has been diverted from its original purpose to capture bone vibrations by placing it in an adjustable 3D-printed enclosure adapted to the participants’ zygomatic bones.
In addition to sensor specifications, Table II presents two signal-to-noise ratio (SNR) measures for each sensor. The first, signal-to-self-noise, is computed as the difference between the $L_{50}^{silence}$ and $L_{50}^{speech}$ values, which will be defined in Section III-E, across the entire speechless-clean split of the dataset, as defined in Section III-D. Notably, all sensors exhibit a high signal-to-self-noise ratio, except for the temple vibration sensor, which is primarily designed to capture much higher amplitude signals. The second measure is the signal-to-external-noise ratio, calculated under an ambient noise level of 90 dB. Since the reference signal is unavailable in this case, we derived this ratio by decomposing it into the signal-to-self-noise minus the external-noise-to-self-noise ratio. This was estimated using the speechless-noisy and speechless-clean subsets, as defined in Section III-D. This decomposition assumes that speakers produce speech at the same loudness in both quiet and noisy environments. However, in practice, the Lombard effect causes speakers to increase their vocal effort in noisy conditions, meaning the reported signal-to-external-noise ratios represent a lower bound on the actual values. Despite this approximation, the signal-to-external-noise ratio remains a direct indicator of each sensor’s noise resilience. Our results show that sensors directly capturing skin vibrations exhibit the highest noise resilience. However, these sensors also tend to have the smallest bandwidth, highlighting the trade-off between noise robustness and frequency range.